Roger's age is 1 1/3 times Steven's age. Eight years ago, Roger's age was twice Steven's age. Find Roger's age?
Amy is twice as old as Jenny. Five years ago, Amy was three times as old as Jenny. Find Jenny's age?

Answer by checkley77(12844)

Roger's age now=4x/3
Steven's age now=x
Eight years age Roger's age was (4x/3)-8
Eight years age Steven's age was x-8
Roger's age 8 years age was twice Steven's age.
(4x/3-8)=2(x-8)
4x/3-8=2x-16
(4x-3*8)/3=2x-16
4x-24=3(2x-16)
4x-24=6x-48
4x-6x=-48+24
-2x=-24
x=-24/-1
x=12 years old is Steven's age now.
4*12/3=48/3=16 years old is Roger's age now.
Proof:
(16-8)=2(12-8)
8=2*4
8=8
-------------------------------------
Amy's age=2x.
Jenny's age =x
Five years ago or
(2x-5)=3(x-5)
2x-5=3x-15
2x-3x=-15+5
-x=-10
x=10 Jenny's age now.
Amy's age now=2*10=20.
Proof:
20-5=3(10-5)
15=3*5
15=15